D (-1,2,3) and E (4,0,-1) . how would i find the angle vector DE [which i found to be (5,-2,4)] makes with the xaxis ?
Answer by Alan3354(69443)

D (-1,2,3) and E (4,0,-1) . how would i find the angle vector DE [which i found to be (5,-2,4)] makes with the xaxis ?
------------------------------
Assuming that F (5,-2,4) is a vector originating at the Origin,
Assign a vector along the x-axis, say G (5,0,0) and find the dot product.
FdotG = |F|*|G|cos
|F| = (25 + 4 + 16)^(1/2) = 3sqrt(5)
|G| = 5
FdotG = 5*5 = 25 (y and z are 0 for G)
cos = 25/15sqrt(5) = 0.745356
Angle = 41.81 degs
----------------------
That's the angle between the 2 vectors. I don't think F is the result of D and E, tho.